OpenStudy (anonymous):

I'm having trouble with getting the answer to this equation once i get the answer i'm supposed to round to the nearest hundreth the equation is: 7.23x+16.51=47.89-2.55x

OpenStudy (lgbasallote):

try multiplying everything by 100 to make everything easier 723x + 1659 = 4789 - 255x does that help?

OpenStudy (shane_b):

@CailynChanel: Can you elaborate on what you've done and what part of it is confusing you?

OpenStudy (anonymous):

the answer is 3.2 the answer is 3.17611336 so if you round the 7 to the 1 it gets you: 3.2

hero (hero):

Even if that is the right answer, you only rounded it to the nearest tenth @CubanLover

OpenStudy (anonymous):

ughhh damn thts right hld on

OpenStudy (lgbasallote):

also it could be a random answer...i for one could have said 4.57291 and no one would have noticed

OpenStudy (anonymous):

but i did my calculations

OpenStudy (lgbasallote):

i dont see any..

OpenStudy (anonymous):

on paper

OpenStudy (lgbasallote):

none of us here have no way of verifying if you're right

OpenStudy (lgbasallote):

so as far as we're concerned..that was a random answer ^_^

OpenStudy (anonymous):

hld on thn ill show u

OpenStudy (lgbasallote):

you should have done that in the first place lol

OpenStudy (shane_b):

http://openstudy.com/code-of-conduct :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

well I subtracted47.89 from each side and got 40.66x+16.51=2.55 then i added 16.51to each side and got40.66x=19.05 then i divided each side by 40.66 and got .469, this is the way I learned how to these problems, but when i go back to check my answer it doesn't come out right.

OpenStudy (anonymous):

here is my work plp

hero (hero):

723x + 1659 = 4789 - 255x Add 225x to both sides; subtract 1659 from both sides: 723x + 255x = 4789 - 1659 978x = 3130 Divide both sides by 978: x = \(\frac{3130}{978}\) Simplify that to decimals..

hero (hero):

I'm pretty sure 9.88 is incorrect

hero (hero):

@CubanLover, you added 7.23 + 2.55 and got 9.88. How?

OpenStudy (lgbasallote):

@CailynChanel if you subtract something from the left or right side of the equation you should do the same to the other side too

OpenStudy (lgbasallote):

7.23x+16.51=47.89-2.55x if you subtract 47.89 from the right side to cancel...you must subtract 47.89 on the left side as well 7.23x + 16.51 - 47.89 = -2.55x does that make sense?

OpenStudy (anonymous):

Wait nevermind i think i have the answer did anybody get 3.22 let me check it

OpenStudy (anonymous):

nope how did u get tht and @Igbasallote did u see my work i posted

OpenStudy (anonymous):

I used a graph

OpenStudy (anonymous):

okkk

OpenStudy (anonymous):

@hero i messed up

OpenStudy (anonymous):

Okay thanks for the help, but that was the right answer

OpenStudy (anonymous):

u welmmmmm ever though we didnt give u the right answer :)